1)     A consumer has a utility function given by

ln U = 5 ln x1 + 3 ln x2

if the budget constraint is given by

10x1 + 14x2 = 124, find

i)                  the optimal quantities of the two goods that the consumer should purchase in order to maximise utility, subject to the budget constraint.

ii)                the value of the consumer’s marginal utility of money at the optimum

iii)              the marginal rate of substitution (MRS) of x1 for x2 and determine its direction at the optimum


 


1
Expert's answer
2021-07-26T23:55:01-0400

a. Solving ln U = 5 ln x1 + 3 ln x2

U=X15 + X23

Maximizing U=X15 + X23

Subject to 10X1+14X2=124

Introduce Lagrangian function

L= X15 + X23 + ƛ(124-10X1-14X2)

Differentiating L with respect to X1, X2 and ƛ

Differentiating L w.r.t. X1

1. = 5X14-10ƛ =0

Differentiating L w.r.t. X2

2. =3X22 - 14 ƛ = 0

Differentiating L w.r.t. ƛ

3. =124-10X1-14X2 =0


Rearranging 1 and 2

Ux/Uy = Px/Py

5X14 / 3X22 = 10/14

X14/X22 = 3/7

X1 = 3

X2=7

b. Consumer's marginal utility

MUx= differentiating utility function w.r.t. to X1

MUx= 5X14

MUy= differentiating utility function w.r.t. to X2

MUy= 3X22


c. MRS x,y = MUx / MUy

MRS x,y = 5X14 / 3X22

This is decreasing as x1 increases and x2 decreases, implying Decreasing MRS. Thus this is well-behaved.